OpenStudy (anonymous):

Simplify quantity x plus 3 over 9 all over quantity x plus 12 over 6

OpenStudy (anonymous):

\[\frac{ \frac{ x+3 }{ 9 } }{ \frac{ x+12 }{ 6 } }=\frac{ x+3 }{ 9 }\times \frac{ 6 }{ x+12}\]